OpenStudy (anonymous):

Choose the missing step in the given solution to the inequality -x - 3 < 13 + 3x.

OpenStudy (anonymous):

-x - 3 < 13 + 3x -4x < 16 x > -4

OpenStudy (anonymous):

what step comes after the first

OpenStudy (anonymous):

-x - 3x < 13 + 3

OpenStudy (anonymous):

-4x - 3 > 13 2x - 3 < 18 -4x - 3 < 13 2x - 3 > 18 these are the answer choies

OpenStudy (anonymous):

all your doing is getting the x's on one side and everything else on the other side

OpenStudy (anonymous):

-4x - 3 < 13

OpenStudy (anonymous):

thank you next question Determine the type of boundary line and shading for the graph of the inequality -4x - y > -6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin. @kelliegirl33

OpenStudy (anonymous):

I am not good at graphing these, except on a number line...sorry

OpenStudy (anonymous):

Choose the correct description of the graph of the inequality -3x > 15.

OpenStudy (anonymous):

Open circle on -5, shading to the left. Closed circle on 5, shading to the left. Open circle on 5, shading to the right. Closed circle on -5, shading to the right.

OpenStudy (anonymous):

-3x > 15 x < -5 (change the sign when dividing by a negative) closed circle on -5, shading to the left

OpenStudy (anonymous):

Which of the following ordered pairs is not a solution to the inequality y < -x + 6? (-3, 0) (-1, 0) (4, 3) (3, -2) --------------------------------------------------------------------------------

OpenStudy (anonymous):

I can tell you about the question before, it will be a dashed line

OpenStudy (anonymous):

y < -x + 6 (3,0) -- (this is correct) (-1,0) -- (this is correct) (4,3) -- NOT CORRECT (3,-2) -- (correct) he one that won't work is (4,3)

OpenStudy (anonymous):

typo...the

OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 2 > -4 and 3x less than or greater to 15 A number line with an open circle on -2, shading to the left, and a closed circle on 5, shading to the right. A number line with a closed circle on -2, shading to the left, and an open circle on 5, shading to the right. A number line with an open circle on -2, a closed circle on 5, and shading in between. A number line with a closed circle on -2, an open circle on 5, and shading in between

OpenStudy (anonymous):

x - 2 > -4 3x <= 15 x > -4 + 2 x <= 5 x > -2 we have x >-2 and x <= 5 a number line with an open circle on -2, closed circle on 5, shading in between

OpenStudy (anonymous):

I have to go now....I hope I helped some :) Bye for now

OpenStudy (anonymous):

Choose the correct description of the graph of the inequality -x greater than or equal to 3.

OpenStudy (anonymous):

Open circle on -3, shading to the left. Closed circle on -3, shading to the left. Open circle on 3, shading to the right. Closed circle on 3, shading to the right

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

-x >= 3 x <= -3 closed circle on -3, shaded to the left

OpenStudy (anonymous):

] Which of the following is not a way to represent the solution of the inequality 3(2x − 1) greater than or equal to 4(2x − 3) − 3? x less than or greater to 6 6 greater than or equal to x A number line with a closed circle on 6 and shading to the left. A number line with a closed circle on 6 and shading to the right

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x − 1 < −12 or x + 12 > 10 x < −11 or x > −2 A number line with an open circle on −11, shading to the left, and an open circle on −2, shading to the right. A number line with an open circle on −11, an open circle on −2, and shading in between. x > −2 or x < −11

OpenStudy (anonymous):

3(2x - 1) >= 4(2x - 3) - 3 6x - 3 >= 8x - 12 - 3 6x - 8x >= -15 + 3 -2x >= - 12 (change signs when dividing by a negative) x <= 6 a number line with closed circle on x, shading to the right is incorrect

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x − 1 < −12 or x + 12 > 10 x < −11 or x > −2 A number line with an open circle on −11, shading to the left, and an open circle on −2, shading to the right. A number line with an open circle on −11, an open circle on −2, and shading in between. x > −2 or x < −11

OpenStudy (anonymous):

x - 1 < -12 x + 12 > 10 x < -12 + 1 x > 10 - 12 x < - 11 x > -2 a number line with open circle on -11,open circle on -2, shading in between is not a solution

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 15 less than or greater to 12 and x + 15 greater than or equal to 5 12 greater than or equal to x + 15 greater than or equal to 5 −3 greater than or equal to x greater than or equal to −10 A number line with a closed circle on −10, a closed circle on −3, and shading in between. A number line with a closed circle on −10, shading to the left, and a closed circle on −3, shading to the right.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

@kelliegirl33 please this is the last one i need

OpenStudy (anonymous):

x + 15 <= 12 x + 15 >= 5 x <= 12 - 15 x >= 5 - 15 x <= -3 x > -10 -10 <= x <= -3 a number line with a closed circle on -10, shading to the left, and a closed circle on -3,shaded to the right is not a solution

OpenStudy (anonymous):

Which of the following is not an equivalent form of the compound inequality x + 15 less than or greater to 12 and x + 15 greater than or equal to 5 12 greater than or equal to x + 15 greater than or equal to 5 −3 greater than or equal to x greater than or equal to −10 A number line with a closed circle on −10, a closed circle on −3, and shading in between. A number line with a closed circle on −10, shading to the left, and a closed circle on −3, shading to the right.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

just did this one

OpenStudy (anonymous):

Determine the type of boundary line and shading for the graph of the inequality −4x − y > −6 Dashed line with shading on the side that includes the origin. Solid line with shading on the side that does not include the origin. Dashed line with shading on the side that does not include the origin. Solid line with shading on the side that includes the origin.

OpenStudy (anonymous):

soory**** that one! this is the last oneeeee

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

I do not know how to do the ones with these graphs....I can tell you that is will be a dashed line. Sorry, that is all I know

OpenStudy (anonymous):

typo ...it

OpenStudy (anonymous):

thank u sso much:)

OpenStudy (anonymous):

no problem :)